3/4

An investor looks at todays yield to maturities in the Wall Street Journal for debt with 10 year maturities. He observes theANNUAL CUPON
Answer #1

3]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 4.49% (YTM of bonds with AAA rating)

nper = 10 (Years remaining until maturity with 1 coupon payment each year)

pmt = 100 * 7.875% (annual coupon payment = face value * coupon rate)

fv = 100 (face value receivable on maturity)

PV is calculated to be $126.80

| Al x fc =PV(4.49%,10,100*7.875%, 100) B C D E F 1 ($126.80)!

Price of bond as % of par =  $126.80 / $100 = 126.80%